Tech Staff Meeting Notes March 30, 2015

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

The fish fry will on May 8 and cost $6.

We discussed the faculty meeting and SRI. The Dean's recent presentation is here

The noise study at GFM revealed the new site is unsuitable (too much road noise). So I guess GFM gets an ISIS4 upgraded instead of a digital one.

We're getting 14 s-13's from South Carolina.

James has rtpd and rtpm running on rhydra3. Mitch will modify achilles to send to there instead of hector. rhydra4 is awaiting orders.

Chris is working on a noise reduction scheme for the maclab.


freewave serial reset installed at 9 of 11 broadbands.


(this is the same list as in the February notes)
We have not had a significant event so pulled the data from here: /gaia/smeagol/dat2/Working/NM/201502/20150224_050505_MAN
Misc Problem
EPRM-dropping out
SJBM - dropping out
FPST- dropping out-spikes
PPLM-dropping out-spikes
SFTN - see image
CVTN - see image The noise at CVTN is due to newly installed monstrous condensor units for the HVAC

Two ISIS3 upgrades remain: ARPT, and the PLAR replacement.

ASAR is closed. This permit was a verbal so Dave will call to find out if we need to formalize the closing. Otherwise we can just go pick up the hardware.

The Basalt is ready for installation at CUET.

NODE Tasks

The win7 replacement for mudworm will be installed this week. Its a 32 bit installation so we don't need to upgrade memory. So they suffer from the Unix Millenium Bug and their epoch times will reset to 1901 in 2038:


James completed link tests with wildworm. Plan to complete the reconfig from Richardson to Wilder by the end of April.




The waveserver as occasional periods with gaps most likely due to missed messages. Steve will investigate.

AZTN/HDBR: Shahram submitted a proposal to TDOT to replace all the sensors with Episensors (including new cabinets, etc). We asked USGS to provide new radios so we can have a continuous buffer of all channels at AZTN. We'll then allow access to NSMP so they can extract event data.


Need to review and update the power systems in 218.

Currently status channels are only available at their respective nodes. Jim asked to have them available on a Memphis waveserver. Mitch will look into it; we need to make sure their scnl's and earthworm logos are unique.


John is helping Gordana Vlahovic get NCCU back online. She has a computer that needs to be configured. Its best to use earthworm exports rather than scream to scream. Steve will help get this done.

CCNC is not going through fgtn and instead repeats through ASTN. The 950MHz shot would work so it could be converted to digital.

We're still repeating CPRT through GTTN. The plan was to put a cell modem at CPRT but John found a potential internet host in the local Library.

Hooray, we finally got the new model to upgrade the Athens DSL.


GMG.HNZ see image
CCRT.EHE see image
CVVA - to be decommissioned?
Mt Gibbs Node
Misc Problem
BCRT.HN* - spikes

We seem to have excessive RT147 problems. Need to follow up with ASL. The FDL committeee agreed to obsolesce the K2's. Jim took possession and had them removed from inventory in about 500 milliseconds.

STAR Node We may have a bad intrusion alarm. Should put a better lock on the cabinet.

Lighting at CCRT has consumed two refteks, an S-13, and several other other pieces of hardware despite having the best grounding and surge supression of any station in the network. We're going to cut our losses and downgrade to ISIS4. The reftek and episensor will find a new home, possibly at CCNC.

The unused pair of ubiquiti radios will be install at Star Mountain. All CERI data will then go over the Ubiquiti link and the EMA can use the Aironet link to transmit the camera data in real-time continuous.


Internet, accessibility, and high noise inducing RF requires we move the MGNC central receive. John presented several possibilities and investigated both the PSN digitizer and the Menlo Park system. We determined that eventually we'll need to convert all nodes to the Menlo Park system so that we can obsolesce XP. But for now, we need to reconfigure North Carolina as in this figure. Better figure. This includes:
Decommission RBNC DONE
Upgrade WSNC to broadband or S-13/Reftek and cell modem. Done.
Rotate CCNC to repeat through fgtn and add to MTTN node. Done.
Establish new node at High Peak.
Transmit MGNC to High Peak.
Repeat TRYN through MGNC.
Transmit GFM to High Peak.
Transmit TVNC to High Peak.
Transmit SMNC to High Peak.

Node hardware at BRBC is pulled. Waiting on tower work at Morgantown. John volunteered to remove the old hardware and install the new, but the tower should be repainted.


CPRT is waiting on county for permission to use Library internet; will probably be ok.

GMG needs a new vertical S-13.

PMTN needs a fence?


Jim provided an update to the spreadsheet detailing the antenna upgrades and ethernet isolation progress at each station. Presumably Green means done and yellow means still needs work.

CJTR is still offline. Waiting for contractors to complete work at the site.

AG Network

See comments in the News and Highlights section on power upgrades.


The committee provided the completed report that was discussed at the February faculty meeting. Chuck would like to see a well equipped, well orgranized and well used FDL. All that takes is time and money of which we have neither. Outside of JerMing's class and Steve Horton's portables, the equipment isn't used much. The FDL wiki is here:


NSMP Station List.


The seisnet group on umdrive is at You will need your uuid and password.